# Yesterday’s News: Benchmarking Multi-Dimensional Out-of-Distribution Generalisation of Misinformation Detection Models

### v1

@@ 9.1 Dataset Access and Licensing
    As mentioned, despite initial releasing the datasets under a CC0 1.0888https://creativecommons.org/publicdomain/zero/1.0/deed.en, the authors have de-accessioned their original repository since January 1st, 2024. Upon request, the authors note their desire to restrict usage to non-commercial research.
    While we want to respect the wishes of our colleagues, we deem the needs of the research community to outweigh these concerns; without access to misinformation texts it is not possible to develop methods for mitigating their effect.
    Despite this, we (plan to) release our dataset under the more restrictive CC BY-NC-SA 4.0999https://creativecommons.org/licenses/by-nc-sa/4.0/deed.en licence. This allows for redistribution and adaption as necessary for academic research, while preventing commercial use-cases and requiring adaptions to maintain these restrictions.
    Furthermore, we (plan to) release the dataset through two media that allow for restricted access. Specifically, we (plan to) use Harvard’s Dataverse (which implements per-file access restrictions) and HuggingFace’s Dataset Hub (which implements repository-level gating). We initially plan to review access applications manually, limiting use-cases to academic research only.
    9.2 Content
    This dataset inherently contains texts that are toxic, hateful, or otherwise harmful to society if disseminated. The dataset itself or any derivative formats of it, like language models, should not be released for non-research purposes.
@@ 9.2 Content
    We have deliberately removed all social media content, and all hyperlinks to such content. We consider such content Personally identifiable information (PII), with limited use in misinformation classification beyond author profiling. Such applications are fraught with ethical problems Mishra et al. (2021), and likely only induce overfitting in text-based classification.
    Appendix A Additional Information on Generalisation Axes and Splits
    1. 1.
      Uniform: for this generalisation axis, we employ the industry standard method of stratified and shuffled train/validation/test splits. We maintain the ratios 80/10/20, respectively, throughout. This is meant to form the baseline set of values, both in terms of ID-OoD articles performance, but also the expected delta between those.
    2. 2.
      Time: misinformation changes quickly, but models have been found to be brittle to articles outside the time-span of the training data Bozarth and Budak (2020); Horne et al. (2020b). We train models on a dataset from a single year, and evaluate performance on all other years. To avoid conflating ‘Time’ and ‘Publisher’ results, we only evaluate on articles from publishers seen in the training set. The proportion of publishers–and as a result, the reliable/unreliable proportions–does change drastically across dataset years.
    3. 3.
      Event: we define events as occurrences that span news articles with a definite time-span and sudden occurrence. Novel events spawn articles with a markedly different vocabulary, introducing words, names, or terms not yet available to the model. Prior work has shown models suffer when evaluated on unseen events Lee et al. (2021); Ding et al. (2022). The dataset contains a multitude of such events, but we focus specifically on the COVID-19 pandemic; an event that particularly notable for its rapid onset and the volume of misinformation it spawned Bradd (2024). Models are trained on a subset of articles not containing COVID-19 keywords, and evaluated on ones that do. The article discovery-process is outlined in more detail in Appendix A.1.
    4. 4.
      Topic: we define topics to be relatively static collections of events, wherein the style or publishers’ opinions change little across years or decades. We discover such topics automatically, bottom-up. Specifically, we apply a heavily modified variant of the BERTopicGrootendorst (2022) algorithm to discover the many thousands of events that occur in each year of the dataset. Each event is textually represented by a TF-IDF vector. To group events into topics, we apply a second round of spectral clustering on the adjacency matrix induced by the inter-event cosine distance matrix. The number of articles contained by a topic differs substantially (see Table 7). We reserve the $n$ smallest topics, which collectively contain roughly 20% of all articles for testing, and train on articles from all other topics.
    5. 5.
      Publisher: all scraped articles are annotated with the news website, outlet, or publisher that produced the article (see Fig. 4). In general, news publishers have some editorial bias or stance, making their corpus of articles distinctly different from that of other publishers. Prior work has found misinformation classifiers to be particularly sensitive to changes in publisher Barrón-Cedeño et al. (2019); Bozarth and Budak (2020). We specifically test for this by reserving the $n$ smallest publishers, which collectively contain roughly 20% of all articles for the test set, and train on articles from all other publishers. We stratify this splitting across political bias, to ensure each political bias is represented equally in both the train and test sets.
    6. 6.
      Political Bias: separate from editorial bias, publishers tend to exhibit a political bias as well. MBFC has annotated all publishers on a left-right political bias continuum. We map all ‘Extreme Left’ and ‘Left’ publishers to a coarser ‘Left’ group, and vice versa for the ‘Right’-side of the political spectrum. All other political biases (‘Center-Left’, ‘Center-Right’, ‘Least Biased’ and ‘Pro-Science’) were mapped to ‘Center’ bias. We reserve either the ‘Left’ or ‘Right’ group for evaluation, and train on the collection of articles from the opposite political bias along with all ‘Center’ biased publishers.
    7. 7.
      Misinformation Type: misinformation presents in various forms. MBFC divides unreliable publishers in three categories: ‘Questionable-Source’ (publishers that exhibit extreme bias or propaganda), ‘Conspiracy-Pseudoscience’ (publisher aligning with known conspiracies or pseudo-scientific practises) and ‘Satire’ (publishers whose content is purposefully false for comedic effect). To test this form of generalisation, we reserve either ‘Questionable-Source’ or ‘Conspiracy-Pseudoscience’ labels for evaluation, and train on all other articles. The test set includes a stratified sample of reliable sources as well.
    A.1 Identifying COVID-19 Articles
    Figure 3: The top figure shows the proportion of articles published that day which discuss the COVID-19 pandemic. The solid line provides a LOESS smoothed trend line. The bottom figure provides a Google search interest for topics that fall under Coronavirus disease 2019. The different years (2020, 2021, 2022) are displayed as banded vertical columns.
    For ‘Event’ splitting, we focus on a single event present over the latter 3 dataset years: the COVID-19 pandemic. We first derive a set of COVID-19 keywords (e.g., ‘sars-cov-2’, ‘lockdown’, ‘mask’), and combine it with the set of keywords defined by Gruppi et al. (2021). We include all articles from 2020-2022 that include any of these terms in the held-out test set. While this is bound to induce a large number of false positives, it ensures no COVID-19 related terms contaminate the models’ learned vocabulary. In Fig. 3 we display the correlation between the found number of articles and the amount of Google search volume.
    A.2 Test Set Correlations
    Year IoU IoU Over Random Test Set Sizes
    Publisher Topic
    2017 0.1183 9.07% 83633 82688
    2018 0.1227 5.62% 352385 375596
    2019 0.1210 6.80% 459787 483084
    2020 0.1386 24.24% 815825 813176
    2021 0.1317 24.82% 843197 832201
    2022 0.1318 32.93% 821905 807449
    Table 6: The amount of article_id overlap between the ‘Publisher’ and ‘Topic’ test sets. Column ‘IoU’ gives the intersection over the union between the two held-out test sets. Column ‘Over Random’ compares this to the achieved value for 10000 simulated draws from a uniform distribution of the same sizes as the actual test set sizes (given in the last two columns).
    Publishers have a natural tendency to prioritise different topics. This is especially prevalent for less mainstream publishers, which have far fewer resources to spend on the breadth of their reporting. It is therefore plausible that the ‘Publisher’ and ‘Topic’ splits have some overlap.
    We test for this by comparing the amount of article level overlap in the held-out test sets of the ‘Publisher’ and ‘Topic’ to random draws of the same size from uncorrelated uniform distributions. The larger this overlap, the more correlated the two test sets are. We display the results of our test in Tab. 6. While we find more overlap than random draws, the effect is not large. Even in the most egregious case, there is only a $0.04$ difference in absolute IoU. As such, we believe the two different test sets measure distinct generalisation aspects.
    Statistic 2017 2018 2019 2020 2021 2022
    Article Counts 0.10M 0.46M 0.59M 1.02M 1.04M 0.99M
    Label Proportion Reliable 61.39% 72.82% 75.91% 70.73% 71.36% 72.00%
    Unreliable 38.61% 27.18% 24.09% 29.27% 28.64% 28.00%
    Publishers Reliable 43 99 134 189 183 184
    Unreliable 49 60 100 249 219 201
    Quantiles 299, 488, 512 266, 448, 512 358, 512, 512 369, 512, 512 366, 512, 512 383, 512, 512
    Event Clustering # Events 2674 6288 7718 7931 8758 8336
    Mean Event Size 38.67 73.38 76.87 128.49 118.24 119.29
    Smallest Topic Size 2.7K 20.5K 24.6K 27.7K 51.3K 48.4K
    Largest Topic Size 38.3K 94.1K 113.1K 192.3K 211.4K 209.4K
    Reliable Labels Left 17.91% 13.32% 11.13% 9.01% 8.76% 9.88%
    Left-Center 24.09% 29.62% 35.19% 36.50% 34.53% 35.01%
    Least Biased 0.93% 2.18% 4.38% 4.91% 5.13% 3.97%
    Right-Center 4.34% 12.66% 9.81% 8.70% 10.38% 9.52%
    Right 14.12% 15.05% 15.41% 11.40% 12.03% 13.08%
    Pro-Science 0.21% 0.53% 0.55%
    Unreliable Labels Satire 1.70% 0.99% 0.82% 0.66% 0.48% 0.42%
    Questionable Source 27.83% 20.14% 17.34% 19.41% 19.00% 18.45%
    Consp.-Pseudosci. 9.09% 6.05% 5.93% 9.20% 9.17% 9.13%
    Table 7: Various statistics on the dataset and labels. Article Counts provides the total number of articles in each dataset year. Label Proportion provides the relative occurrence of reliable vs. unreliable articles, whereas Publishers provides the number of such publishers present. Section Truncated Token Counts provides the mean and the 25, 50, 75th quantiles of the number of tokens per article. Note that is after truncation at 512; the raw articles tend to be much longer. Event Clustering provides the number and average size (in articles) of events. We also provide the number of events belonging to the smallest and largest topics. Sections Reliable Labels and Unreliable Labels provides the proportion of each MBFC label in the dataset, split into their class label.
    Appendix B Dataset Processing & Statistics
@@
    Table 8: Size of the datasets, in terms of millions of articles, after each step of cleaning, per year. The lower percentage gives the step-to-step reduction in size. The ’Total’ column computes reduction relative to ’Original’.
    We downloaded the initial NELA corpora from the Harvard Dataverse [^10] The corpora have since been de-accessioned, and can longer be downloaded. We expand on this in Sec. 9, under a CC0 1.0 licence [^11] https://creativecommons.org/publicdomain/zero/1.0/deed.en.
@@
    Table 9: Publishers removed for being news aggregation sites, with article counts post de-aggregation.
@@
    Table 10: Publishers removed for being social media sites, with article counts post de-aggregation.
    B.1 Re-Labelling
    We manually re-mapped all URL domains to a set of publishers consistent across years, excluding all news aggregation platforms and social media sites (see Tables B and B).
    It should be noted that the 2018 edition of NELA did not contain URLs, making relabelling considerably more difficult.
@@ B.1 Re-Labelling
    Previews To read the full blog, please check out the complete post on <source>
@@
    Table 11: Frequent duplicated articles.
@@
    The article must have a mean token length greater than \${VALUE} characters 2
    The article must have a mean token length less than \${VALUE} characters 10
@@
    Table 12: Filtering rules.
    B.2 Deduplication
    1. 1.
      we only keep the first (as determined by publication date) instance of exact duplicates for each publisher. Besides plagiarism, the most common duplicates are due to errors in scraping or parsing, resulting in artefacts (see Table 11)
  • 2.
    we remove all articles with duplicate titles or URLs, across the entire dataset. This primarily removes articles that are updated at a later stage (e.g., live-blogs or summaries), or result from URL re-directions 3. we remove all articles if more than 5% of its sentences are duplicates from the same publisher
    The former two (‘Exact Deduplication’ in Tab. 8) can handle document-level duplicates, like plagiarised articles or updated blog-posts, but miss near duplicates. The latter should filter out near duplicates, like lightly edited documents, (‘Near Deduplication’ in Tab. 8).
    An added benefit of these deduplication steps is the identification of common scraping errors. In fact, we found this to be the most frequent form of duplication. We list some common texts in Table 11. Since the majority of such errors were automated responses unique to individual publishers, we could filter these out with relative ease using the described ‘Exact Deduplication’ approaches.
  • B.3 Cleaning & Filtering
    We normalize all punctuation and remove any embedded URLs, HTML or Markdown markup. SpaCy’s Montani et al. (2023) en_core_web_sm was used to annotate all tokens in the corpora. We identify self-references as named entities with a large longest common substring relative to the publisher’s name. We also mask any sentences which SpaCy flags as being part of an email, URL or twitter handle. Finally, we standardize the NELA copyright poisoning, applying it to all dataset years equally.
@@ B.2 Deduplication
    To ensure all articles contained enough grammatical text to reasonably classify, we removed any article which did not abide by the rules delineated in Table 12. As a final step, we use Lingua Stahl (2024) to filter out any non-English texts.
@@ B.4 Event & Topic Clustering
    To generate the metadata necessary for the ‘Topic’ generalisation form (see Sec. 4 or App. A), we opted for a bottom-up approach. This involved first clustering the dataset into thousands of fine-grained events, before clustering the event clusters into overarching topics.
@@
    Figure 4: Schema for the generated metadata database. ’PK’ indicates primary key, ’FK’ denotes a foreign key relationship.
    After embedding an entire year of the dataset, we first reduce their dimensionality using mini-batched PCA, and whiten the data. We construct a vocabulary over all the documents, by aggregating mini-batches of vocabularies. We then apply UMAP dimensionality reduction, using the PCA solution as initialisation, and the cosine distance as the distance metric. Clustering was performed via HDBScan to mini-batches of the embeddings, assigning all articles to their nearest event cluster. Using the dataset-wide vocabulary, we generate a single TF-IDF representation matrix.
    After mini-batching, we generate a hierarchical clustering on top of the event-based TF-IDF matrix. We merge any events with a distance below 0.8 to mitigate the effect of mini-batching, and re-construct the TF-IDF representations. We deem these the final event representations. Table 7 provides the total number of events and the average number of articles contained within each.
    Finally, we apply a round of spectral clustering to these representations, compressing these events into 10 groups, which we deem topics. The numbers of events in each topic varies (see Table 7).
@@ B.5 Metadata
    Article content is stored in arrow files, accessed using HuggingFace’s datasets library HuggingFace (2021). We store all generated metadata in a duckdb database Mühleisen and Raasveldt (2024). Fig. 4 depicts the metadata schema. Table ’Sources’ provides information sourced from MBFC on different publishers, whereas tables ’Articles’ and ’Topics’ provides information on the produced articles and events/topics, respectively. The articles are all given unique article identifiers. This ensures we can quickly generate relevant splits of the dataset and link these to article-level predictions.
@@ B.6 Publisher-level Metadata
    All publisher metadata is stored in the ’Sources’ table. Each source is identified by a unique name, linked in a 1-to-many relationship to articles in the ’Articles’ table. The url column provides a link to the MBFC page with the source metadata. The label column provides the MBFC label, also shown in the lower rows of Table 7. In column bias, the political bias (one of ’Extreme Left’, ’Left’, ’Left-Center’, ’Least Biased’, ’Pro-Science’, ’Right-Center’, ’Right’ or ’Extreme Right’) of the publisher is provided. The factuality column provides an ordinal value for the propensity of a publisher to report factual news. Columns country and type provide information about the country of origin of the publisher, and their primary form of publication (e.g., TV, blog posts), respectively. In column popularity, the average number of visits to the publisher’s main website is provided as an ordinal categorical value. credibility provides an overall assessment of the publishers’ credibility, aggregating all other variables into a single score. Finally, the conspiracy and pseudosci columns provide the ‘strength’ of the conspiracy or pseudo-science source. The larger this value is, the more these sources deviate from the public or scientific consensus. This is only provided for publishers labelled as ‘Conspiracy-Pseudoscience’.
    Figure 5: The authorship of articles in the dataset, aggregated into deciles. The right most column of each panel represents the top 10% most prolific publishers, the left most the bottom 90%.

### v2

@@ 10.1 Dataset Access and Licensing
We aim to make misinfo-general as easy to use as possible, but have had to make some restrictions. The dataset contains texts that are toxic, hateful, or otherwise harmful to society if disseminated. The dataset itself or any derivative formats of it, like language models, should not be released for non-research purposes.
The NELA corpora were initially released under a CC0 1.0101010https://creativecommons.org/publicdomain/zero/1.0/deed.en, essentially being released to the public domain. From January 1st, 2024, the NELA authors have de-accessioned their repository. Upon request, the authors note their desire to restrict usage to non-commercial research.
Given the potentially harmful content, and our colleagues wishes, we (re-)release our dataset under a more restrictive CC BY-NC-SA 4.0111111https://creativecommons.org/licenses/by-nc-sa/4.0/deed.en license. This allows for redistribution and adaption as necessary for academic research, while preventing commercial use-cases and requiring adaptions to maintain these restrictions.
@@
We have deliberately removed all social media content, and all hyperlinks to such content. We consider such content Personally identifiable information (PII), with limited use in misinformation classification beyond author profiling. Such applications are fraught with ethical problems Mishra, Yannakoudakis, and Shutova (2021), and likely only induce overfitting in text-based classification.
We have released misinfo-general through two media that allow for restricted access. Specifically, we use Harvard’s Dataverse (which implements per-file access restrictions) and HuggingFace’s Dataset Hub (which implements repository-level gating). We plan to review access applications manually, limiting use-cases to academic research only.
Appendix A Misinformation Dataset Survey
{longtblr}
[ theme = misinfodatasets, caption = A long table with an (inexhaustive) sampling of misinformation datasets. Each row provides a single dataset, with name and citation, along with the labelling granularity (see Sec. 3), the dataset size (where units ‘k’ and ‘M’ denote thousands and millions, respectively), and a short description of how the dataset authors generated misinformation labels., label = tab:misinfo_datasets, notea = MechanicalTurk: crowdsourced lay volunteers, noteb = Weibo Community Management Center: credible Weibo users can report posts, notec = Buzzfeed: a digital media company, noted = Snopes: expert journalist website for debunking misinformation, notee = PolitiFact: expert journalist website for fact checking politicians, notef = GossipCop: a defunct website dedicated to fact-checking celebrity rumors, noteg = HealthNewsReviews: a defunct website dedicated to reviewing medical claims, noteh = Poynter: a global, non-profit organization with annotations from partnered organizations, notei = Google Fact Check Tool: a unified API for fact-checking annotations ] colspec = >p0.250 >p0.100 >p0.050 X , width = 1.0rowhead = 1, rowfoot = 0, hlines = 0.25pt, solid, hline1, 2, Z = 1pt, solid, Dataset & Label Size Description
[…]
MediaEval15 Boididou et al. Claim 12k The authors generated a list of events which were verified as true or false, and a collection of tweets discussing these events. The tweets were manually verified
[…]
TSHP-17 Rashkin et al. (2017) Publisher 70k Trusted news articles were sampled from the Gigaword News corpus, whereas unreliable news was sampled from specific publishers.
Kaggle Fake News Risdal (2016) Publisher 13k The authors scraped articles from unreliable sources using a third-party tool. No reliable articles were included
[…]
PHEME Zubiaga, Liakata, and Procter (2017) Claim 5.8k Tweets related to 5 mainstream events were manually annotated as unverified rumour or verified
[…]
Weibo17 Jin et al. (2017) Article, Publisher 10k They take posts reported as false from trusted users, and take articles from mainstream publishers for their factual class
Some Like it Hoax Tacchini et al. (2017) Publisher 15.5k Articles were scraped from Facebook groups dedicated to sharing scientific or pseudo-scientific articles
Fake vs Satire Golbeck et al. (2018) Publisher 0.5k Articles were sampled from identified satire or fake news sites. The authors constrained the number of articles per publisher to ensure a diverse publisher set. All ambiguous cases were removed
[…]
FakeNewsCorpus Pathak and Srihari (2019) Article, Publisher 0.7k ~700 articles were taken from questionable source publishers, and used as misinformation, and 26 expert labelled factual news articles. Satire and unverifiable news were explicitly excluded.
QProp Barrón-Cedeño et al. (2019) Publisher 51k Uses MBFC to assign articles the label of their publisher. They manage to sample from 104 different sources, although only include 10 progandistic sources.
[…]
WeChat Wang et al. (2020) Article 4k The authors collected articles flagged by WeChat users. A small subset was annotated by experts, while a larger subset was unannotated, meant for unsupervised training
CoAID Cui and Lee (2020) Claim, Article, Publisher 3.7k Misinformation articles about the COVID19 pandemic were scraped directly from various fact-checking sources. Factual articles were scraped from 9 reliable publishers. Claims were scraped from official government sites
[…]
MCFEND Li et al. (2024) Article 24K Articles annotated by various fact-checking organisations around the world were collected, and manually mapped to a single annotation schema.
Appendix B misinfo-general Processing & Statistics
@@ Appendix A Misinformation Dataset Survey
Table B.1: Size of the datasets, in terms of millions of articles, after each step of cleaning, per year. The lower percentage gives the step-to-step reduction in size. The ’Total’ column computes reduction relative to ’Original’.
We downloaded the initial NELA corpora from the Harvard Dataverse, under a CC0 1.0 license [^12] https://creativecommons.org/publicdomain/zero/1.0/deed.en. The corpora have since been de-accessioned, and can longer be downloaded. We expand on this in Section 10.1.
B.1 Re-Labelling
We manually re-mapped all URL domains to a set of publishers consistent across years, excluding all news aggregation platforms and social media sites (see Tables B.4 and B.4). It should be noted that the 2018 edition of NELA did not contain URLs, making relabelling in this manner impossible.
Table B.2: Publishers removed for being news aggregation sites, with article counts post de-aggregation.
@@ B.1 Re-Labelling
Table B.3: Publishers removed for being social media sites, with article counts post de-aggregation.
@@
Table B.4: Frequent duplicated articles.
@@
Previews To read the full blog, please check out the complete post …
@@
B.2 Deduplication
  1. 1.
    we only keep the first (as determined by publication date) instance of exact duplicates for each publisher. Besides plagiarism, the most common duplicates are due to errors in scraping or parsing, resulting in artifacts (see Table B.4)
  2. 2.
    we remove all articles with duplicate titles or URLs, across the entire dataset. This primarily removes articles that are updated at a later stage (e.g., live-blogs or summaries), or result from URL re-directions
  3. 3.
    we remove all articles if more than 5% of its sentences are duplicates from the same publisher
The former two (‘Exact Deduplication’ in Table B.1) can handle document-level duplicates, like plagiarized articles or updated blog-posts, but miss near duplicates. The latter should filter out near duplicates, like lightly edited documents, (‘Near Deduplication’ in Table B.1).
An added benefit of these deduplication steps is the identification of common scraping errors. In fact, we found this to be the most frequent form of duplication. We list some common texts in Table B.4. Since the majority of such errors were automated responses unique to individual publishers, we could filter these out with relative ease using the described ‘Exact Deduplication’ approaches.
B.3 Cleaning & Filtering
We normalize all punctuation and remove any embedded URLs, HTML, or Markdown markup. SpaCy’s Montani et al. (2023) en_core_web_sm was used to annotate all tokens in the corpora. We identify self-references as named entities with a large longest common substring relative to the publisher’s name. We also mask any sentences which SpaCy flags as being part of an email, URL or twitter handle. Finally, we standardize the NELA copyright poisoning, applying it to all dataset years equally.
Table B.5: Filtering rules.
@@ B.3 Cleaning & Filtering
The article must have a mean token length greater than \${VALUE} 2
The article must have a mean token length less than \${VALUE} 10
@@
To ensure all articles contained enough grammatical text to reasonably classify, we removed any article which did not abide by the rules delineated in Table B.5. As a final step, we use Lingua Stahl (2024) to filter out any non-English texts.
@@ B.4 Event & Topic Clustering
To generate the metadata necessary for the ‘Topic’ generalization form (see Section 5 or Appendix C), we opted for a bottom-up approach. This involved first clustering the dataset into thousands of fine-grained events, before clustering the event clusters into overarching topics.
@@
After embedding an entire year of the dataset, we first reduce their dimensionality using mini-batched PCA, and whiten the data. We construct a vocabulary over all the documents, by aggregating mini-batches of vocabularies. We then apply UMAP dimensionality reduction, using the PCA solution as initialization, and the cosine distance as the distance metric. Clustering was performed via HDBScan to mini-batches of the embeddings, assigning all articles to their nearest event cluster. Using the dataset-wide vocabulary, we generate a single TF-IDF representation matrix.
After mini-batching, we generate a hierarchical clustering on top of the event-based TF-IDF matrix. We merge any events with a distance below 0.8 to mitigate the effect of mini-batching, and re-construct the TF-IDF representations. We deem these the final event representations. Table B.6 provides the total number of events and the average number of articles contained within each.
Finally, we apply a round of spectral clustering to these representations, compressing these events into 10 groups, which we deem topics. The numbers of events in each topic varies (see Table B.6).
Figure B.1: Schema for the generated metadata database. ’PK’ indicates primary key, ’FK’ denotes a foreign key relationship.
@@ B.5 Metadata
Article content is stored in arrow files, accessed using HuggingFace’s datasets library HuggingFace (2021). We store all generated metadata in a duckdb database Mühleisen and Raasveldt (2024). Figure B.1 depicts the metadata schema. Table ’Sources’ provides information sourced from MBFC on different publishers, whereas tables ’Articles’ and ’Topics’ provides information on the produced articles and events/topics, respectively. The articles are all given unique article identifiers. This ensures we can quickly generate relevant splits of the dataset and link these to article-level predictions.
@@ B.6 Publisher-level Metadata
All publisher metadata is stored in the ’Sources’ table. Each source is identified by a unique name, linked in a 1-to-many relationship to articles in the ’Articles’ table. The url column provides a link to the MBFC page with the source metadata. The label column provides the MBFC label, also shown in the lower rows of Table B.6. In column bias, the political bias (one of ’Extreme Left’, ’Left’, ’Left-Center’, ’Least Biased’, ’Pro-Science’, ’Right-Center’, ’Right’ or ’Extreme Right’) of the publisher is provided. The factuality column provides an ordinal value for the propensity of a publisher to report factual news. Columns country and type provide information about the country of origin of the publisher, and their primary form of publication (e.g., TV, blog posts), respectively. In column popularity, the average number of visits to the publisher’s main website is provided as an ordinal categorical value. credibility provides an overall assessment of the publishers’ credibility, aggregating all other variables into a single score. Finally, the conspiracy and pseudosci columns provide the ‘strength’ of the conspiracy or pseudo-science source. The larger this value is, the more these sources deviate from the public or scientific consensus. This is only provided for publishers labelled as ‘Conspiracy-Pseudoscience’.
B.7 Article Statistics
Table B.6: Various statistics on the dataset and labels. Article Counts provides the total number of articles in each dataset year. Label Proportion provides the relative occurrence of reliable vs. unreliable articles, whereas Publishers provides the number of such publishers present. Section Truncated Token Counts provides the mean and the 25, 50, 75th quantiles of the number of tokens per article. Note that is after truncation at 512; the raw articles tend to be much longer. Event Clustering provides the number and average size (in articles) of events. We also provide the number of events belonging to the smallest and largest topics. Sections Reliable Labels and Unreliable Labels provides the proportion of each MBFC label in the dataset, split into publisher categories.
Statistic 2017 2018 2019 2020 2021 2022
Article Counts 0.10M 0.46M 0.59M 1.02M 1.04M 0.99M
Publishers Reliable 43 99 134 189 183 184
Unreliable 49 60 100 249 219 201
Q25 299 266 358 369 366 383
Q50 488 448 512 512 512 512
Q75 512 512 512 512 512 512
Event Clustering # Events 2674 6288 7718 7931 8758 8336
Mean Event Size 38.67 73.38 76.87 128.49 118.24 119.29
Smallest Topic Size 2.7K 20.5K 24.6K 27.7K 51.3K 48.4K
Largest Topic Size 38.3K 94.1K 113.1K 192.3K 211.4K 209.4K
Label Proportion reliable+ 61.39% 72.82% 75.91% 70.73% 71.36% 72.00%
unreliable- 38.61% 27.18% 24.09% 29.27% 28.64% 28.00%
reliable+ Left 17.91% 13.32% 11.13% 9.01% 8.76% 9.88%
Left-Center 24.09% 29.62% 35.19% 36.50% 34.53% 35.01%
Least Biased 0.93% 2.18% 4.38% 4.91% 5.13% 3.97%
Right-Center 4.34% 12.66% 9.81% 8.70% 10.38% 9.52%
Right 14.12% 15.05% 15.41% 11.40% 12.03% 13.08%
Pro-Science 0.21% 0.53% 0.55%
unreliable- Satire 1.70% 0.99% 0.82% 0.66% 0.48% 0.42%
Questionable Source 27.83% 20.14% 17.34% 19.41% 19.00% 18.45%
Consp.-Pseudosci. 9.09% 6.05% 5.93% 9.20% 9.17% 9.13%
We present various statistics at the article level for each iteration of the final misinfo-general dataset in Table B.6. This includes the number of articles, the average number of tokens per article (post truncation), the number of events and the size of topics, and finally the label proportions, both in aggregated form as reliable+ or unreliable- , but also per publisher category. In total, the dataset comprises some 2B tokens once truncated. Without truncation, this is likely substantially higher.
In Table B.7 we present descriptions of the various topics in the 2020 split of misinfo-general. These topics are latent, and automatically generated from the inter-event distance matrix. The space in which the articles were clustered into events is high-dimensional, and each cluster was assumed to be non-convex. As a result, it is difficult to find an ’average’ representation of any event, and even more difficult to find one for a topic (i.e., a cluster of clusters). Instead, we derived topic descriptions by sampling events from each topic, and using the BERTopic generated event representation.
Table B.7: An example of the topics present in the 2020 split of misinfo-general. As described in Section 5, to form the OoD set, the $k$ smallest topics are chosen, such collectively they (approximately) comprise 20% of all articles.